OpenStudy (anonymous):

help plse: solve z-4 / z+5 = 1/7 z=?

OpenStudy (anonymous):

7z-28 = z+5 6z = 33 z = 11/2

OpenStudy (anonymous):

could you show the steps plse

OpenStudy (anonymous):

(z-4)/(z+5) = 1/7 cross multiply 7(z-4) = 1(z+5) 7z - 28 = z+5

OpenStudy (anonymous):

how did you arrive at z=11/2

OpenStudy (anonymous):

see what i did earlier if 6z = 33 divide both sides by 3 2z = 11 z = 11/2

OpenStudy (anonymous):

ok gotcha can you show me the steps to this one plse: add simplify if possible 3/rv^2 + 7/r^2v =

OpenStudy (anonymous):

(3r+7v)/r^2v^2

OpenStudy (anonymous):

can you show me the steps to this one plse: add simplify if possible 3/rv^2 + 7/r^2v =